Country A has a per capita income of $4,000 per year and Country B has a per capita income of S40.000 per year. Assume Country As per capita income grows at a rate of 5% per year and Country Bs per capita income grows at a rate of 290 per year How many times larger is Country Bs income per capita today? 10 times larger How many times larger will Country Bs income per capita be in 70 years? Use the rule of 70 to figure this out. 1.25 times larger 1. a. b.

Need help, please show work
Answer #1

(a) Income per capita, Country B / Income per capita, Country A = $40,000 / $4,000 = 10 times

(b) Using rule of 70, Doubling period = 70 / Annual growth rate.

Doubling period, Country A = 70 / 5 = 14

In 70 years, country A's GDP per capita will double (70/14) = 5 times.

After 70 years, country A's GDP per capita = $4,000 x 2 x 2 x 2 x 2 x 2 = $128,000

Doubling period, Country B = 70 / 2 = 35

In 70 years, country B's GDP per capita will double (70/35) = 2 times.

After 70 years, country B's GDP per capita = $40,000 x 2 x 2 = $160,000

Income per capita, Country B / Income per capita, Country A = $160,000 / $128,000 = 1.25 times